# Safe Model-Based Multi-Agent Mean-Field Reinforcement Learning

### v1

@@ 5 Experiments
Results. The target distribution $\rho_{0}$ in Figure 1(a) has a relatively high entropy due to a wide horizontal spread of $67\%$ of the maximum entropy. To achieve vertical spread, we require additional 18 percentage points of entropy as a safety constraint. Concretely, we use $p=85\%$ to set the hard constraint $C=p\log(k^{2})$ and proceed by optimizing the policy profile in Section 4. Due to the lack of an analytical solution for Section 4, we use a policy profile trained with known dynamics as a benchmark. We show that $\bm{\pi}_{n}^{*}$ converges to this policy profile in $n=80$ episodes. Figure 2(a) shows two phases of the learning process. During the first 60 episodes, the performance is volatile, and then the confidence bound shrinks, i.e., the random exploration reduces. Once the epistemic uncertainty around the true dynamics is tight, the model exploits it rapidly by episode 80.
@@
By comparing Safe-$\text{M}^{3}$-UCRL to the benchmark $\text{M}^{3}$-UCRL introduced in [43], which does not impose any safety constraint, we empirically show that Safe-$\text{M}^{3}$-UCRL satisfies the safety constraint during the entire execution. In contrast, $\text{M}^{3}$-UCRL quickly violates it (see Figure 2(b)). In Figure 2, we use a city map of Shenzhen to show that Safe-$\text{M}^{3}$-UCRL improves service accessibility in low-demand areas. Figure 1(b) shows that $\text{M}^{3}$-UCRL with known dynamics learns how to satisfy the demand $\rho_{0}$ effectively, but with the discussed drawback of violating safety constraint. Figure 1(c) shows that Safe-$\text{M}^{3}$-UCRL with known dynamics improves on that by satisfying safety constraint by distributing taxis to residential areas in the northwest and northeast. Finally, Figure 1(d) emphasizes the capability of Safe-$\text{M}^{3}$-UCRL to learn complex system dynamics while the policy profile $\bm{\pi}_{n}^{*}$ simultaneously converges towards the results achieved with known dynamics with the number of episodes $n$ passed.
@@
In Appendix D, we provide the details on the parameters used during the training and exhaustive performance analysis. The results in this section are generated assuming the infinite regime. At the same time, in Appendix D, we also discuss the performance when the policy profile $\bm{\pi}_{n}^{*}$ is applied to the finite number of individual agents. The code we used to train and evaluate Safe-$\text{M}^{3}$-UCRL is available at https://github.com/mjusup1501/safe-m3-ucrl.
6 Conclusion
@@
We present a novel formulation of a mean-field model-based reinforcement learning problem incorporating safety constraints. Safe-$\text{M}^{3}$-UCRL addresses this problem by leveraging epistemic uncertainty in the unknown transition dynamics model and employing a log-barrier approach to ensure conservative satisfaction of the constraints. Through experimental evaluations, we showcase the effectiveness of our algorithm in matching the demand distribution in a shared mobility service while consistently upholding service accessibility. In the future, we believe that integrating safety considerations in intelligent multi-agent systems will have a crucial impact on various applications, such as firefighting robots and drone/robot search-and-rescue operations in complex and confined spaces (e.g., mining accidents or tunnel collapses).
Acknowledgments and Disclosure of Funding
@@
The authors would like to thank Mojmír Mutný for the fruitful discussions during the course of this work. Matej Jusup acknowledges support from the Swiss National Science Foundation under the research project DADA/181210. This publication was made possible by an ETH AI Center doctoral fellowship to Barna Pasztor. Francesco Corman acknowledges Grant 2023-FS-331 for Research in the area of Public Transport. The authors would also like to thank the Shenzhen Urban Transport Planning Center for collecting and sharing the data used in this work.
@@
Appendix A Relationship Between True and Statistical Mean-Field Distributions
In this section, we describe the procedure for computing constants $C_{n,t}$ (defined in Section 4) at every episode $n=1,\ldots,N$ and step $t=1,\ldots,T$. These are necessary for establishing the connection between the mean-field distributions $\mu_{n,t}$ in the original system with known transition dynamics $f$ (Section 3.1) and the mean-field distributions $\tilde{\mu}_{n,t}$ in the system induced by a calibrated statistical model $\tilde{f}_{n-1}$ (Section 4, Section 3.2 and Assumption 4). In particular, Corollary 1 shows that the safety of the original system $h_{C}(\mu_{n,t})$ is guaranteed if $h_{C}(\tilde{\mu}_{n,t})\geq L_{h}C_{n,t}$ for a safety constant $C_{n,t}\geq W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$ and Lipschitz constant $L_{h}$ introduced in Assumption 6. We use the following result from [43, Lemma 5] to make a connection between the mean-field distributions in the true dynamics and statistical model dynamics:
@@ Appendix A Relationship Between True and Statistical Mean-Field Distributions
$W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})\leq 2\beta_{n-1}\overline{L}_{n-1}^{t-1}\sum_{i=0}^{t-1}\int_{\mathcal{S}}\|\bm{\sigma}_{n-1}(s,\mu_{n,i},\pi_{n,t}(s,\mu_{n,i}))\|_{2}\mu_{n,i}(ds),$
where $\overline{L}_{n-1}=1+2(1+L_{\pi})(L_{f}+2\beta_{n-1}L_{\sigma})$.
@@
for $t=1,\ldots,T$, where the first inequality is due to Lemma 2 while the second one follows since maximum upper bounds expectation. Hence, we can set $C_{n,t}$ as $tK_{n,t}\max_{z\in\mathcal{Z}}\|\bm{\sigma}_{n-1}(z)\|_{2}$, and calculate $\max_{z\in\mathcal{Z}}\|\bm{\sigma}_{n-1}(z)\|_{2}$ once at the beginning of each episode $n$. Notice that our learning protocol (Algorithm 1) does not require computing $C_{n,0}$ at the initial step $t=0$ since the mean-field distributions share the initial state, i.e., $\tilde{\mu}_{n,0}=\mu_{n,0}$. It is worth noting that in certain models, such as Gaussian Processes, this upper bound provides a meaningful interpretation. Specifically, $\bm{\sigma}_{n-1}(z)$ represents the epistemic uncertainty of the model, which tends to decrease monotonically as more data is observed.
@@ Appendix B Examples of Safety Constraints
In this section, we show some important classes of safety constraints $h_{C}(\cdot)$ satisfying Assumption 6.
B.1 Entropic Safety
Entropic safety serves the purpose of controlling the dispersion of the mean-field distribution, which is useful in many applications, such as vehicle repositioning (see Section 5). A natural way of defining entropic safety is via differential entropy, but in general, the differential entropy is not Lipschitz continuous due to the unboundedness of the natural logarithm. Nevertheless, the issue can be easily circumvented by considering $\varepsilon$-smoothed differential entropy $H^{\varepsilon}:\mathcal{P}(\mathcal{S})\rightarrow\mathbb{R}_{+}$. For $\varepsilon>0$ and $C>0$ we define $\varepsilon$-smoothed differential entropy and associated entropic safety constraint $H_{C}^{\varepsilon}(\cdot)$ as:
$H^{\varepsilon}(\mu)$ $=-\int_{\mathcal{S}}\log(\mu(s)+\varepsilon)\mu(ds)$
$H^{\varepsilon}_{C}(\mu)$ $=H^{\varepsilon}(\mu)-C$
To show that $H^{\varepsilon}_{C}(\cdot)$ satisfies Assumption 6 let $h_{C}(\cdot):=H^{\varepsilon}_{C}(\cdot)$ and assume that $\mathcal{S}\subset\mathbb{R}^{p}$ is a compact set. First, note that $f(x)=\log(x+\varepsilon)$ is $\frac{1}{\varepsilon}$-Lipschitz continuous for $\varepsilon>0$, i.e., $\varepsilon f(x)$ is 1-Lipschitz. Second, for every $S\subseteq\mathcal{S}$ and $L$-Lipschitz function $f:\mathcal{S}\rightarrow\mathbb{R}$, a function $g(S)=\int_{\mathcal{S}}f(s)\mu(ds)$ is $L$-Lipschitz due to the boundedness of $f$. The following derivation shows that $H_{C}^{\varepsilon}(\cdot)$ is $\frac{1}{\varepsilon}$-Lipschitz continuous:
@@ B.1 Entropic Safety
$=\left|\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)\mu^{\prime}(ds)-\int_{\mathcal{S}}\log(\mu(s)+\varepsilon)\mu(ds)\right|$
$=\left|\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)(\mu^{\prime}-\mu)(ds)-\int_{\mathcal{S}}\log(\mu(s)+\varepsilon)\mu(ds)+\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)\mu(ds)\right|$
$\leq\left|\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)(\mu^{\prime}-\mu)(ds)\right|+\left|\int_{\mathcal{S}}\log(\mu(s)+\varepsilon)\mu(ds)-\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)\mu(ds)\right|$
$\leq\left|\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)(\mu^{\prime}-\mu)(ds)\right|+\frac{1}{\varepsilon}W_{1}(\mu,\mu^{\prime})$
$\leq M\left|\int_{\mathcal{S}}(\mu^{\prime}-\mu)(ds)\right|+\frac{1}{\varepsilon}W_{1}(\mu,\mu^{\prime})$
$=\underbrace{M(\mu^{\prime}-\mu)(\mathcal{S})}_{\overline{M}}+\frac{1}{\varepsilon}W_{1}(\mu,\mu^{\prime})$
$=\overline{M}+\frac{1}{\varepsilon}W_{1}(\mu,\mu^{\prime}),$
@@
In Section 5 we use the discrete equivalent of entropic safety, i.e., Shannon entropy, for our experiments. [45, Proposition 8] shows that Shannon entropy is Lipschitz continuous with respect to the scaled Wasserstein 1-distance, i.e., with respect to $\frac{1}{n}W_{1}(\cdot)$ known as Ornstein’s distance.
B.2 Distribution Similarity
We can define safety by preventing the mean-field distribution $\mu$ from diverging from a prior distribution $\nu_{0}$ by quantifying the allowed dissimilarity between the two distributions:
$h_{C}(\mu;\nu_{0})$ $=C-D(\mu,\nu_{0}),$
where $C>0$ and $D:\mathcal{P}(\mathcal{S})\times\mathcal{P}(\mathcal{S})\rightarrow\mathbb{R}_{+}$ is the distance function between probability measures. Commonly used distances are Wasserstein p-distance for $p\geq 1$ and $f$-divergences such as KL-divergence, Hellinger distance, and total variation distance.
A concrete example of a distance function $D(\cdot)$ that satisfies Assumption 6 is Wasserstein 1-distance:
$W_{1}^{C}(\mu,\nu_{0})$ $=C-W_{1}(\mu,\nu_{0})$
@@ B.2 Distribution Similarity
[14] shows the triangle inequality of the Wasserstein p-distance for probability measures on separable metric spaces. [17] shows that Wasserstein 1-distance induces a metric space $(\mathcal{P}(\mathcal{S}),W_{1})$ over probability measures. The result now trivially follows from the reverse triangle inequality:
@@
Avoiding risky distributions can be modelled by setting $W_{1}^{C}(\mu,\nu_{0})=W_{1}(\mu,\nu_{0})-C$ with the proof of Assumption 6 being equivalent to the above.
Weighted safety constraints. In applications that require emphasis on certain regions of the state space, we can generalize the above safety constraints by introducing the weight function $w:\mathcal{S}\rightarrow\mathbb{R}_{+}$. We extend Section B.1 to weighted-differential entropy by defining $H^{w,\varepsilon}(\mu):=-\int_{\mathcal{S}}w(s)\log(\mu(s)+\varepsilon)\mu(ds)$ and Section B.2 by considering weighted Wasserstein 1-distance $W_{1}^{w}(\mu,\nu_{0})=\sup_{f:\text{Lip}(f)\leq 1}\int_{\mathcal{S}}w(s)f(s)(\mu-\nu_{0})(ds)$. Here, we use Kantorovic-Rubinstein dual definition of Wasserstein 1-distance $W_{1}(\mu,\nu_{0})=\sup_{f:\text{Lip}(f)\leq 1}\int_{\mathcal{S}}f(s)(\mu-\nu_{0})(ds)$, where $f:\mathcal{S}\rightarrow\mathbb{R}$ is a continuous function and $\text{Lip}(f)$ denotes the minimal Lipschitz constant for $f$.

### v2

@@ 5.2. Vehicle Repositioning Problem
Results. The entropy of the target distribution, $\rho_{0}$ in Figure 2(a), already achieves $p=0.67$ of the maximum due to a wide horizontal spread. To achieve vertical spread, we require an additional 18 percentage points of entropy as a safety constraint. Concretely, we use $p=0.85$ to set the threshold as the proportion of maximum entropy and proceed by optimizing the policy profile in Section 4. Due to the lack of an analytical solution for Section 4, we use a policy profile trained under known transitions as a benchmark. We observe that the learned policy profile $\bm{\pi}_{n}^{*}$ converges to the policy profile under known transitions in $n=80$ episodes. Figure 1(d) shows two phases of the learning process. During the first 60 episodes, the performance is volatile, but once the epistemic uncertainty around true transitions is tight, the model exploits it rapidly by episode 80.
@@
In Figure 1(f), we empirically show that Safe-M3-UCRL satisfies safety constraints during the entire execution. In Figure 3, we use a city map of Shenzhen to show that Safe-M3-UCRL improves service accessibility in low-demand areas. Figure 2(b) shows that M3-UCRL under known transitions learns how to satisfy the demand $\rho_{0}$ effectively at the cost of violating safety constraint (see Figure 1(f)). Figure 2(c) shows that Safe-M3-UCRL under known transitions improves safety by distributing vehicles to residential areas in the northwest and northeast. Finally, Figure 2(d) emphasizes the capability of Safe-M3-UCRL to learn complex transitions while the policy profile $\bm{\pi}_{n}^{*}$ simultaneously converges towards the results achieved under known transitions with the number of episodes $n$ passed.
@@
In Section D.3, we provide the details on the parameters used during the training and exhaustive performance analysis. The results in this section are generated assuming the infinite regime. At the same time, in Section D.5, we showcase that in the finite regime the policy profile $\bm{\pi}_{n}^{*}$ can be successfully applied to millions of individual agents in real-time, which might be of particular importance to real-world practitioners. The code we use to train and evaluate Safe-M3-UCRL is available in our GitHub repository Jusup and Janik (2023).
6. Conclusion
@@
We present a novel formulation of the mean-field model-based reinforcement learning problem incorporating safety constraints. Safe-M3-UCRL addresses this problem by leveraging epistemic uncertainty under an unknown transition model and employing a log-barrier approach to ensure conservative satisfaction of the constraints. Beyond the synthetic swarm motion experiment, we showcase the potential of our algorithm for real-world applications by effectively matching the demand distribution in a shared mobility service while consistently upholding service accessibility. In the future, we believe that integrating safety considerations in intelligent multi-agent systems will have a crucial impact on various applications, such as autonomous ride-hailing, firefighting robots and drone/robot search-and-rescue operations in complex and confined spaces.
{acks}
@@
The authors would like to thank Mojmír Mutný for the fruitful discussions during the course of this work. Matej Jusup acknowledges support from the Swiss National Science Foundation under the research project DADA/181210. This publication was made possible by an ETH AI Center doctoral fellowship to Barna Pasztor. Francesco Corman acknowledges Grant 2023-FS-331 for Research in the area of Public Transport. Andreas Krause acknowledges that this research was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement no. 815943 and the Swiss National Science Foundation under NCCR Automation, grant agreement 51NF40 180545. Ilija Bogunovic acknowledges that the work was supported in part by the EPSRC New Investigator Award EP/X03917X/1. The authors would also like to thank the Shenzhen Urban Transport Planning Center for collecting and sharing the data used in this work.
@@
Appendix A Relationship Between Mean-Field Distributions under True and Estimated Transitions
In this section, we describe the procedure for computing constant $C_{n,t}$ (defined in Section 4) at every episode $n{}={}1,\ldots,N$ and step $t{}={}1,\ldots,T$. This is necessary for establishing the connection between the mean-field distribution $\mu_{n,t}$ in the original system under known transitions $f$ (see Section 3.1) and the mean-field distribution $\tilde{\mu}_{n,t}$ in the system induced by a calibrated statistical model $\tilde{f}_{n-1}$ (see Section 4, Section 3.3 and Assumption 4). In particular, Corollary 2 shows that the safety in the original system $h_{C}(\mu_{n,t})$ is guaranteed with high probability if $h_{C}(\tilde{\mu}_{n,t}){}\geq{}L_{h}C_{n,t}$ for a safety constant $C_{n,t}{}\geq{}W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$ and Lipschitz constant $L_{h}$ introduced in Assumption 6. We use the following result from (Pásztor et al., 2023, Lemma 5) to make a connection between the mean-field distributions under true and estimated transitions.
@@ Appendix A Relationship Between Mean-Field Distributions under True and Estimated Transitions
$W_{1}(\tilde{\mu}_{n,t},\mu_{n,t}){}\leq{}2\beta_{n-1}\overline{L}_{n-1}^{t-1}\sum_{i=0}^{t-1}I_{n,t},$
where $I_{n,t}{}={}\int_{\mathcal{S}}\|\bm{\sigma}_{n-1}(s,\mu_{n,i},\pi_{n,t}(s,\mu_{n,i}))\|_{2}\mu_{n,i}(ds)$ and $\overline{L}_{n-1}{}={}1{}+{}2(1{}+{}L_{\pi})(L_{f}{}+{}2\beta_{n-1}L_{\sigma})$.
@@
for $t=1,\ldots,T$, where the first inequality is due to Lemma 4 while the second one follows since maximum upper bounds expectation. Hence, we can set $C_{n,t}$ as $tK_{n,t}\max_{z\in\mathcal{Z}}\|\bm{\sigma}_{n-1}(z)\|_{2}$, and calculate $\max_{z\in\mathcal{Z}}\|\bm{\sigma}_{n-1}(z)\|_{2}$ once at the beginning of each episode $n$. Notice that our learning protocol (see Algorithm 1) does not require computing $C_{n,0}$ at the initial step $t=0$ since the mean-field distributions share the initial state, i.e., $\tilde{\mu}_{n,0}=\mu_{n,0}$. It is worth noting that in certain models, such as Gaussian Processes, this upper bound provides a meaningful interpretation. Specifically, $\bm{\sigma}_{n-1}(z)$ represents the epistemic uncertainty of the model, which tends to decrease monotonically as more data is observed.
@@ Appendix B Examples of Safety Constraints
In this section, we show some important classes of safety constraints $h_{C}(\cdot)\geq 0$ satisfying Assumption 6.
B.1. Entropic Safety
Entropic safety serves the purpose of controlling the dispersion of the mean-field distribution, which is useful in many applications, such as vehicle repositioning (see Section 5). A natural way of defining entropic safety is via differential entropy, but in general, the differential entropy is not Lipschitz continuous due to the unboundedness of the natural logarithm. Nevertheless, the issue can be easily circumvented by considering $\varepsilon$-smoothed differential entropy $H^{\varepsilon}:\mathcal{P}(\mathcal{S})\rightarrow\mathbb{R}_{\geq 0}$. For $\varepsilon>0$ and $C\geq 0$ we define $\varepsilon$-smoothed differential entropy and associated entropic safety constraint $H_{C}^{\varepsilon}(\cdot)$ as
$H^{\varepsilon}(\mu)$ ${}:={}-\int_{\mathcal{S}}\log(\mu(s){}+{}\varepsilon)\mu(ds)$
$H^{\varepsilon}_{C}(\mu)$ ${}:={}H^{\varepsilon}(\mu)-C$
To show that $H^{\varepsilon}_{C}(\cdot)$ satisfies Assumption 6 let $h_{C}(\cdot){}:={}H^{\varepsilon}_{C}(\cdot)$ and assume that $\mathcal{S}\subset\mathbb{R}^{p}$ is a compact set. First, note that $f(x){}={}\log(x{}+{}\varepsilon)$ is $\frac{1}{\varepsilon}$-Lipschitz continuous for $\varepsilon>0$, i.e., $\varepsilon f(x)$ is 1-Lipschitz. Second, for every $S\subseteq\mathcal{S}$ and $L$-Lipschitz function $f:\mathcal{S}\rightarrow\mathbb{R}$, a function $g(S)=\int_{\mathcal{S}}f(s)\mu(ds)$ is $L$-Lipschitz due to the boundedness of $f$. The following derivation shows that $H_{C}^{\varepsilon}(\cdot)$ is $\frac{1}{\varepsilon}$-Lipschitz continuous.
@@ B.1. Entropic Safety
[…]
$\begin{split}={}&\left|\int_{\mathcal{S}}\log(\mu^{\prime}(s){}+{}\varepsilon)(\mu^{\prime}-\mu)(ds)-\int_{\mathcal{S}}\log(\mu(s){}+{}\varepsilon)\mu(ds)\right.\\ &\left.{}+{}\int_{\mathcal{S}}\log(\mu^{\prime}(s){}+{}\varepsilon)\mu(ds)\right|\end{split}$
$\begin{split}{}\leq{}{}&\left|\int_{\mathcal{S}}\log(\mu^{\prime}(s){}+{}\varepsilon)(\mu^{\prime}-\mu)(ds)\right|\\ &{}+{}\left|\int_{\mathcal{S}}\log(\mu(s){}+{}\varepsilon)\mu(ds)-\int_{\mathcal{S}}\log(\mu^{\prime}(s){}+{}\varepsilon)\mu(ds)\right|\end{split}$
[…]
@@
In Section 5, Appendix D and Appendix E, we use the discrete equivalent of entropic safety, i.e., Shannon entropy, for our experiments because of the discrete representation of the mean-field distribution. (Polyanskiy and Wu, 2016, Proposition 8) shows that Shannon entropy is Lipschitz continuous with respect to the scaled Wasserstein 1-distance, i.e., with respect to $\frac{1}{n}W_{1}(\cdot)$ known as Ornstein’s distance.
B.2. Distribution Similarity
We can define safety by preventing the mean-field distribution $\mu$ from diverging from a prior distribution $\nu_{0}$ by quantifying the allowed dissimilarity between the two distributions
$h_{C}(\mu;\nu_{0})$ ${}:={}C-D(\mu,\nu_{0}),$
where $C\geq 0$ and $D:\mathcal{P}(\mathcal{S})\times\mathcal{P}(\mathcal{S})\rightarrow\mathbb{R}_{\geq 0}$ is the distance function between probability measures. Commonly used distances are Wasserstein p-distance for $p{}\geq{}1$ and $f$-divergences such as KL-divergence, Hellinger distance, and total variation distance.
A concrete example of a distance function $D(\cdot)$ that satisfies Assumption 6 is Wasserstein 1-distance
$W_{1}^{C}(\mu,\nu_{0})$ ${}:={}C-W_{1}(\mu,\nu_{0})$
@@ B.2. Distribution Similarity
Clement and Desch (2008) shows the triangle inequality of Wasserstein p-distance for probability measures on separable metric spaces. Dudley (2018) shows that Wasserstein 1-distance induces a metric space $(\mathcal{P}(\mathcal{S}),W_{1})$ over probability measures. The result now trivially follows from the reverse triangle inequality
@@
Avoiding risky distributions can be modeled by setting $W_{1}^{C}(\mu,\nu_{0}){}={}W_{1}(\mu,\nu_{0})-C$ with the proof of Assumption 6 being equivalent to the above.
Weighted safety constraints. In applications that require emphasis on certain regions of the state space, we can generalize the above safety constraints by introducing the weight function $w:\mathcal{S}\rightarrow\mathbb{R}_{\geq 0}$. We extend Section B.1 to weighted-differential entropy by defining
$H^{w,\varepsilon}(\mu){}:={}-\int_{\mathcal{S}}w(s)\log(\mu(s){}+{}\varepsilon)\mu(ds)$
and Section B.2 by considering weighted Wasserstein 1-distance
$W_{1}^{w}(\mu,\nu_{0}){}:={}\sup_{f:\text{Lip}(f){}\leq{}1}\int_{\mathcal{S}}w(s)f(s)(\mu-\nu_{0})(ds).$
Here, we use Kantorovic-Rubinstein dual definition of Wasserstein 1-distance
$W_{1}(\mu,\nu_{0}){}:={}\sup_{f:\text{Lip}(f){}\leq{}1}\int_{\mathcal{S}}f(s)(\mu-\nu_{0})(ds),$
where $f:\mathcal{S}\rightarrow\mathbb{R}$ is a continuous function and $\text{Lip}(f)$ denotes the minimal Lipschitz constant for $f$.